OpenStudy (anonymous):

Algebra 1 help anyone?

jagr2713 (jagr2713):

whats ur question?

OpenStudy (anonymous):

Jose has$2.15 in nickels and dimes. He has ten more nickles than dimes. How many of each does he have?

OpenStudy (anonymous):

And I need to show my work, so please don't just give the answer. :)

OpenStudy (anonymous):

I know that n-d=10 and n(0.1)+d(0.05)=245

OpenStudy (anonymous):

sorry, not used to nickels and dimes :P

jagr2713 (jagr2713):

is there an answer choice

OpenStudy (anonymous):

no. but I do know that the answer is 13 dimes and 23 nickles, i just need to show my work, and i don't know how to solve it.

OpenStudy (anonymous):

@lilsis76 and @jagr2713 are u there?

jagr2713 (jagr2713):

yea just thinking

OpenStudy (anonymous):

ok :)

jagr2713 (jagr2713):

how did u get 13dimes

OpenStudy (lilsis76):

im thinking too lol sorry, kinda sick :)

OpenStudy (anonymous):

Oh, my teacher gave me the answer, and I'm supposed to show how to solve it @jagr2713 and thats too bad :( @lilsis76

OpenStudy (lilsis76):

it doesnt say how many of the nickels or dimes there are?

OpenStudy (anonymous):

nope, but I guess there must be 36 :)

OpenStudy (lilsis76):

lol okay let me see

OpenStudy (anonymous):

I think that the next step is 0.05(245-0.1n)=245 but I'm not sure

undeadknight26 (undeadknight26):

I BELIEVE IN YOU @eyust707 and @lilsis76! GET EM!

OpenStudy (anonymous):

@undeadknight26 do you know if the equation above is correct? ^

OpenStudy (lilsis76):

haha sorry doing this slowly blah . dumb cold! lol Jose has 2.15 in N & D. he has 10 more N than D How many of each does he have? is this the quesion, Right?

undeadknight26 (undeadknight26):

:O I have no involvement yet ;p Arent you just seeing how many dimes and nickles make up $2.15? O.o

OpenStudy (anonymous):

no @undeadknight26 there are 10 more nickels than dimes and yes @lilsis76

undeadknight26 (undeadknight26):

Oh makes sense...This takes too much energy and im 2 lazy sowwy </3

OpenStudy (lilsis76):

okay im trying it out, and i am on this equations .05x + .10y = 2.15

OpenStudy (lilsis76):

well i think the equation is right lol gotta see now

OpenStudy (eyust707):

Yes then just write one more equation...

OpenStudy (anonymous):

ok so (im usind n and d for nickels and dimes) so there is only one unknown variable, that would be 0.05(245-0.1d)=245

OpenStudy (anonymous):

is that correct do you know?

OpenStudy (lilsis76):

okay got it. are you ready @Lucinda12345 ?

OpenStudy (lilsis76):

so i gave you the equation: .05x+.10y=2.15 where the x is nickels and y is dimes so we'll change it to n and d

OpenStudy (anonymous):

yes :)

OpenStudy (lilsis76):

.05n + .10d =2.15

OpenStudy (lilsis76):

now your going to want to mult. by 100 to get rid of the decimals so now we have 5n + 10d = 215

OpenStudy (lilsis76):

just remember x is the nickels and so we have to do (n-5) for d (or y) so, 5n + 10 (n-5) = 215 from there you now get 5n+10 (n-5)=215

OpenStudy (anonymous):

ok

OpenStudy (lilsis76):

from there: you can do the math to get: 5n+10n-50=215 simplify 15n-50=215 simplify more 15n=265 simplify even more divide both sides by 15 there are 17 nickels. NOW! fill 17 for the beginning equation. and tell me what you get for Dimes.

OpenStudy (anonymous):

oh you know that it is $2.45 not $2.15

OpenStudy (lilsis76):

what? haha well then just do everything the same exact way and tell me what you get. the way Im explaining it is the way I like things to be explained to me. so lets work it out together and tell me how you are solving it along the way. let me check my answers and the the problem too

OpenStudy (anonymous):

ok so if it is 2.45 them umm... hold on

OpenStudy (lilsis76):

lol okay, ya i know im still doing it. no calc. so using laptop

OpenStudy (anonymous):

So that would equal 13 dimes, and 23 nickles! Thank you for the help :)

OpenStudy (lilsis76):

:) BEST response please. and Ya, its best to talk it out. right?

OpenStudy (anonymous):

yup i gave you best response a while ago :) and I also fanned you

OpenStudy (lilsis76):

im looking up at the to i lost my spot. what did the back of the book say the answer was? lol

OpenStudy (anonymous):

13 and 23

OpenStudy (lilsis76):

lol found it lol

OpenStudy (lilsis76):

okay your answer checks out. Have fun with the Homework. I got physics >:/ BOOOOOO lol